The set of values of a for which (a-1)x2-(a+1)x+a-1>=0 is true for all x>=2

Grade:11

1 Answers

Yashit Gupta
6 years ago
(a-1)x^2-(a+1)x+a-1>=0.....(1)x=2Substituting the value of x in (1) eq.It becomes-(a-1)4-(a+1)2+a-1>=04a-4-2a-2+a-1>=03a-7>=0a>=7/3==》a is greater than or equal to zeroSo the value of a lies between 7/3 and infinity.So,the required set is (7/3,infinity)7